OpenStudy (anonymous):

a particle moving on the x axis has position x(t)= 2t^3+3t^2-36t+40 find the total distance traveled by the particle during the first 3 seconds The answer is 61 keep getting |x(1)-x(0)|+|x(3)-x(1)|=35 I got the zeros by factoring the derivative of x(t) (v(t)).

OpenStudy (unklerhaukus):

\[x(t)= 2t^3+3t^2-36t+40\] \[=\int\limits_0^3 2t^3+3t^2-36t+40 \text d t\]\[=\left.\frac{t^4}{2}+\frac 3 2t^3-18t^2+40t\right|_0^3\]\[=\frac{(3)^4}{2}+\frac 3 2(3)^3-18(3)^2+40(3)\] \[=\frac {81}{2}+\frac 3 2 \times 27 -18\times 9 +120\]

OpenStudy (unklerhaukus):

except that doesn't give 61, there is probably a few error in my working

OpenStudy (anonymous):

This is in the related rate section I can't use that lol

OpenStudy (anonymous):

I'll give you a medal anyways :)

OpenStudy (unklerhaukus):

the second them the 3/2 bit should have been 3/3=1

OpenStudy (anonymous):

I have to find the zero's of the velocity because that is when the "particle" changes direction then calculate the distance

OpenStudy (anonymous):

zeros*

OpenStudy (unklerhaukus):

oh, so to find the zeroes take the first derivative of x(t)

OpenStudy (anonymous):

yeah but I did that I just keep getting 35? idk what I did wrong

OpenStudy (unklerhaukus):

\[x(t)= 2t^3+3t^2-36t+40 \] \[\dot x=\frac{\text d x }{\text d t}=6t^2+6-36=6(t^2-5)\]

OpenStudy (unklerhaukus):

looks like it is turning at t= √5

OpenStudy (anonymous):

6t*

OpenStudy (anonymous):

not 6 lol

OpenStudy (unklerhaukus):

?

OpenStudy (anonymous):

x'(t)= 6t^2+6t-36

OpenStudy (unklerhaukus):

oh yeh

OpenStudy (anonymous):

hehe

OpenStudy (unklerhaukus):

\[x(t)=2t^3+3t^2−36t+40\] \[x'=6t^2+6t−36=6(t^2+t-6)=6(t-2)(t+3)\]

OpenStudy (unklerhaukus):

so turning at t= 2 and t= 3

OpenStudy (anonymous):

thats what I did wrong! thanks!

OpenStudy (unklerhaukus):

t=-3 i mean

OpenStudy (unklerhaukus):

so you should break the integral into \[\int_0^2x(t)\text d t+\int_2^3x(t)\text d t\]